OpenStudy (anonymous):

The perimeter of a rectangle is 34yd, and the area of the rectangle is 64yd^2 . Find the dimensions of the rectangle. Can anyone help me please.

OpenStudy (anonymous):

\[P=2(x+y)\]

OpenStudy (anonymous):

\[2(x+y)=34\] \[A=xy=64\] solve simultaneously

OpenStudy (anonymous):

\[xy=64\] \[x+y=17\]

OpenStudy (anonymous):

can you solve simultaneously

OpenStudy (anonymous):

how did you get the 17?

OpenStudy (anonymous):

2 divided into the 34 right

OpenStudy (anonymous):

w=16 and L=18?

OpenStudy (anonymous):

yes 34/2=17

OpenStudy (anonymous):

it is 60yd^2 not 64yd^2